Solve the equation for R
7 = 2 (r+1) exponent2 -3

Answer by oscargut(2103)

Hi, the answer is: r = sqrt(5)-1 or r = -sqrt(5)-1
If you need more help or details you can contact me at:
mthman@gmail.com
Thanks